Sam holds a lit match in front of a lens (f = 25.0
Answer #1

(a) real ans upside down

(b) 1/f = 1/v - 1/u

1/25 = 1/v + 1/37.5

image distance v = 75 cm

(c) m = h2/h1 = v/u

h2 = 1.5*75/37.5 = 3 cm

h2 = 3 cm